# FRAMES: Boosting LLMs with A Four-Quadrant Multi-Stage Pretraining Strategy

### v1

@@ 3.1 Settings
For 3B models, our pretraining data is sourced from various domains, including books Gao et al. (2020), blogs Baumgartner et al. (2020), patents Sharma et al. (2019), Common Crawl Penedo et al. (2024), and Wikipedia. It comprises 500B tokens each for Chinese and English, similar to the Matrix dataset Zhang et al. (2024a). For 1.3B models, we randomly select 100B tokens from the SlimPajama dataset Soboleva et al. (2023).
@@
We test on 3B parameter model for the main experiments, with a batch size of 640 and a context window length of 8192. The Adam optimizer is used for training within the Megatron framework. In addition we also validate on a 1.3B model, comparing it with other models based on PPL, PD, and QuRating metrics. More details could be found in Appendix B.1.
@@
We use benchmarks including MMLU Hendrycks et al. (2020), CMMLU Li et al. (2023), BBH Suzgun et al. (2022), and CEVAL Huang et al. (2023) for 3B models, covering tasks that assess multi-domain knowledge and complex reasoning. For 1.3B models, we evaluate the models’ performance on the following benchmarks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020), which include tasks such as knowledge question answering and commonsense reasoning. We apply in-context learning and select examples based on task characteristics. Standard accuracy serves as the final metric for all tasks.
@@ 3.2 Main Results
Table 4 shows our experimental results. FRAMES significantly outperforms the Random baseline, with a 9.0% improvement in average performance. For specific tasks, it achieves notable gains of 15.3% on MMLU and 18.2% on CMMLU. In the 1.3B model, FRAMES also surpasses all single-dimension baselines, with 4.1% improvement in average, and a maximum of 6.4% gain in a single task. These findings validate the effectiveness of FRAMES.
Figure 1 and 5 shows the model’s benchmark performance over training steps. The model remains stable initially, with a significant performance increase after 90K steps. Figure 6 illustrates the training loss, showing four declines that align with benchmark performance improvements. This confirms the strong link between training loss and benchmark performance: when dataset phases are equally divided, lower training loss results in better performance. Additionally, FRAMES effectively enhances the model’s emergent capabilities, allowing it to acquire foundational skills quickly and better understand data patterns in later stages, improving overall performance. Through the loss smoothness analysis in Appendix B.6, we found that the loss curve of FRAMES has the lowest high-frequency energy proportion of only 0.02%, significantly lower than Random and PDPC. This indicates that FRAMES can make the model converge more stably and reduce the impact of gradient fluctuations during training.
@@ 3.3 Ablation Study
From the perspective of four quadrants, do the conclusions of Sections 2.1 and 2.3 still hold? We test various quadrant combinations and find that:
@@
More results can be found in Appendix B.7.
@@ 3.4 Analysis
We investigate the characteristics of data within the four quadrants by extracting 1,000 samples from each and analyzing their quality using four raters from QuRating. As shown in Figure 8, the quality distributions are similar across all dimensions for samples from each quadrant, which suggests that model-perceived measures like PPL and PD don’t significantly correlate with human cognition-based measures, such as knowledge, quality, and diversity. FRAMES maintains consistent data quality throughout the training process, preventing the model from focusing solely on low-quality data at any stage.
@@
We extract 230,000 samples from 23 domains within our training dataset, covering areas such as Arxiv, Law, Code, and Math. Using the 1.3B reference model, we calculate the PPL of all samples and present the PPL distributions in Figure 13. Significant differences are observed across various domains, with mean PPL values ranging from 5 to 22 and varying degrees of variance. Data from domains such as Wikipedia typically exhibit lower PPL, whereas data from Reddit show higher PPL. This implies that sorting data by PPL could cause significant shifts in domain representation between initial and later training stages, potentially exposing the model to overly homogeneous data during each training stage. Sachdeva et al. (2024) highlight that data diversity enhances model performance, and an uneven distribution may lead to a decline in performance.
@@
We extract 1 million samples from each of the four quadrant distributions and visualize them, as shown in Figure 9. We observe that each quadrant contains data from a wide range of sources, indicating that FRAMES effectively ensures data diversity throughout the training process. Notably, there are distinct differences among the quadrants. For instance, $Q_{2}$ contains a large number of domain-specific datasets such as Wikipedia, and StackExchange. This aligns with common cognitive understanding, as these datasets may feature more complex sentence structures and terminology, which large models can learn effectively while small models may struggle to master, thus suitable for learning in the later stages of training. This also explains the excellent performance in knowledge-intensive tasks, as the texts in Wikipedia contain numerous factual knowledge. Additionally, the data distribution in $Q_{3}$ is relatively balanced, which is advantageous for the initial training phase. Presenting data from $Q_{3}$ at the beginning of the training process allows the model to learn from sufficiently diverse data, thereby establishing a strong foundation for subsequent learning stages.
@@
To assess distinct semantic properties within each quadrant’s data, we randomly select 1K samples from each quadrant and devise 10 language-text related traits for GPT-4o evaluation. For simplicity and precision, all traits are formulated as yes-or-no questions, and we calculate the percentage of samples meeting each trait’s criteria. Details of all traits are listed in Appendix D. As shown in Figure 10, the data across all four quadrants exhibit similar semantic properties. In addition, we use T5 Raffel et al. (2023) to obtain dense vectors of the samples and perform dimensionality reduction using t-SNE, as shown in Figure 11. We could observe that the data in the four quadrants do not have obvious distinctions at the semantic level. This indicates no obvious difference in the semantic distribution of data between the initial and later training stages, ensuring a diverse range of data is consistently encountered throughout the training process, and avoiding the collapse of the model into a certain preference, thereby damaging generalization performance.
Figure 10: Semantic feature analysis of different quadrants.
@@ 4 Related Works
Multi-stage pretraining has emerged as a pivotal strategy in the development of LLMs, enabling models to better capture and utilize diverse data characteristics by dividing the training process into distinct phasesPavlova (2025); Zhao et al. (2024); Liu et al. (2021b); Tan et al. (2022). Liu et al. propose a multi-stage pretraining method that leverages various granularities of information, significantly boosting model performanceLiu et al. (2021b). Yildiz et al. explore the use of multi-stage prompting to improve translation tasks, demonstrating its effectiveness in enhancing downstream applicationsTan et al. (2022).
In LLM pretraining, data preprocessing is key to ensuring dataset quality. Traditional methods use expert rules to filter low-quality data Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and remove duplicates Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). These approaches enhance quality but may lack semantic depth. To improve semantic selection, strategies involve using targeted sources or proxy models Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024). Classifiers automate selection, like the logistic regression model used by Du et al. (2022), with others employing complex scoring Zhang et al. (2024b); Sachdeva et al. (2024). QuRating Wettig et al. (2024) uses multiple raters for nuanced evaluation.
@@ 5 Conclusion
In this study, we propose the Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES), a novel approach designed to boost the performance of LLMs by systematically organizing the pretraining process into four stages. Guided by the principle of achieving significant loss reductions four times, FRAMES employs a strategic partitioning of pretraining data into four quadrants based on PPL and PD. The experimental results, showing a 9.03% average improvement over random sampling, underscore the effectiveness of FRAMES in optimizing pretraining data organization.
@@ B.1 Experimental settings
We train reference models with 100M and 1.3B parameters on 500B tokens randomly sampled from the collected dataset, utilizing the Llama architecture Touvron et al. (2023). We use the 1.3B model to compute PPL and both the 100M and 1.3B models to calculate PD. For the main experiments, we train a 3B model. We set the batch size to 640 and the context window length to 8192. The initial learning rate is $2\times 10^{-4}$, with a warm-up phase of 375M tokens. We apply cosine learning rate scheduling with a weight decay of 0.1. We use the Adam optimizer to train the model within the Megatron framework Shoeybi et al. (2019). Furthermore, we validate our approach on a 1.3B model. Using the same settings as the 3B model, we also compare it with models trained by sequencing based on the PPL, PD, and QuRating Wettig et al. (2024) metrics.
@@ B.2 Benchmark Accuracy of Two-stage Pretraining Based on PPL
Methods Steps MMLU CMMLU
Random 30K 25.43 25.88
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 26.65 26.19
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 25.29 25.53
Random 60K 25.94 25.84
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.53 25.48
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 33.35 35.37
Random 95K 24.84 25.57
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.96 25.97
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.63 42.63
@@ B.3 Benchmark Accuracy of Two-stage Pretraining Based on PD
Table 6 shows the model’s accuracy as training steps increase under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. The model consistently outperforms the Random model at most steps. Notably, at 95K steps, it exceeds the Random setting by 2.05% on MMLU and 1.63% on CMMLU, which validates PD as an effective metric. This leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
Methods Steps MMLU CMMLU
Random 30K 25.43 25.88
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.65 26.19
Random 60K 25.94 25.84
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.23 25.10
Random 95K 24.84 25.57
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.89 27.20
@@ B.4 S-shape Functions with a=10 and a=35
Figure 13: PPL distribution across different sources.
Figure 14: PD Distribution across different sources.
@@ B.5 Data Distribution
Figure 13 and 14 show the PPL and PD distribution of samples in our training set. The PPL of data from different domains shows significant variations, with mean values ranging from 5 to 22 and varying degrees of variance. Sorting data by PPL may lead to imbalanced domain representation between early and late training stages, exposing the model to overly homogeneous data at each stage and potentially degrading performance. In contrast, the distribution of PD is quite similar across different data sources, ensuring that each training phase includes diverse data, thereby maintaining pretraining efficiency and enhancing model performance.
@@
B.6 Smoother LOSS Curve from FRAMES

### v4

@@ 3.1 Settings
For 3B models, our pretraining data is derived from various domains, including books Gao et al. (2020), blogs Baumgartner et al. (2020), patents Sharma et al. (2019), Common Crawl Penedo et al. (2024), and Wikipedia. It comprises a total of 1T tokens, with 500B tokens each for Chinese and English, akin to the Matrix dataset Zhang et al. (2024a). For 1.3B models, we randomly select 100B tokens from the SlimPajama dataset Soboleva et al. (2023).
@@
We train 100M and 1.3B models on i.i.d subsets of the collected dataset, comprising 500B tokens, to compute PPL and PD. And we test on 3B models for the main experiments, with a batch size of 640 and a context window length of 8192. The Adam optimizer is used for training within the Megatron framework. In addition, we validate our approach on a 1.3B model, employing the same configuration. More details of model structure and training could be found in Appendix B.1.
Baselines
We compare FRAME with random data sequence and PDPC Zhang et al. (2025). In addition we also validate on a 1.3B model, comparing it with other models based on PPL, PD, and QuRating metrics. More details of baselines could be found in Appendix B.1.
@@
For 1.3B models, we evaluate the models’ performance on the following benchmarks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020). For 3B models, we additionally use benchmarks including MMLU Hendrycks et al. (2020), CMMLU Li et al. (2023), BBH Suzgun et al. (2022), and CEVAL Huang et al. (2023), covering more complicated tasks. We apply in-context learning and select examples based on task characteristics. Standard accuracy serves as the final metric for all tasks.
@@ 3.2 Main Results
Table 4 shows our experimental results. FRAME significantly outperforms the Random baseline, with a 6.1% improvement in average performance. For specific tasks, it achieves notable gains of 15.3% on MMLU and 18.2% on CMMLU. For the 1.3B model, as illustrated in Table 3, FRAME also surpasses all the baselines, with 4.1% improvement in average, and a maximum of 6.4% gain in a single task. These findings validate the effectiveness of FRAME. Moreover, in our experiments with the 1.3B and 3B models, distinct pretraining datasets are utilized for each model. Notably, FRAME consistently yield substantial performance gains across all evaluated benchmarks. These results highlight the robustness and generalizability of our method, demonstrating its effectiveness across varying data sources and model architectures.
Figure 1 shows the model’s benchmark performance over training steps. The model remains stable initially, with a significant performance increase after 90K steps. Figure 5 illustrates the training loss, showing four declines that align with benchmark performance improvements. This confirms the strong link between training loss and benchmark performance: when training stages are equally divided, lower training loss results in better performance. Additionally, FRAME effectively enhances the model’s emergent capabilities, allowing it to acquire foundational skills quickly and better understand data patterns in later stages, improving overall performance. Through the loss smoothness analysis in Appendix B.7, we found that the loss curve of FRAME has the lowest high-frequency energy proportion of only 0.02%, significantly lower than Random and PDPC. This indicates that FRAME can make the model converge more stably and reduce the impact of gradient fluctuations during training.
@@ 3.3 Ablation Study
Furthermore, we could observe a slower convergence rate when training on high PD data. This phenomenon can be attributed to the difficulty of high PD samples, which are particularly challenging for the model in the early stage of training when its capacity is still limited. However, as training progresses, the model’s capabilities gradually improve. Since the weak and strong models fit low PD data similarly, even as the model enhances, the training loss will not further decrease, leading to faster convergence. On the other hand, high PD data, with the improvement of model abilities, can further guide the model to learn new features and continue reducing training loss.
@@
From the perspective of four quadrants, do the conclusions of Sections 2.1 and 2.2 still hold? We test various quadrant combinations and find that:
@@
In addition, the $Q_{3}\rightarrow Q_{1}\rightarrow Q_{2}$ setup, where $f_{merge}$ is not applied, shows a performance drop between the second and third stages. This suggests that $f_{merge}$ is necessary. More results can be found in Appendix B.8.
@@ 3.4 Analysis
We investigate the characteristics of data within the four quadrants by extracting 1,000 samples from each and analyzing their quality using four raters from QuRating. As shown in Figure 7, the quality distributions are similar across all dimensions for samples from each quadrant, which suggests that model-perceived measures like PPL and PD don’t significantly correlate with human cognition-based measures, such as knowledge, quality, and diversity. FRAME maintains consistent data quality throughout the training process, preventing the model from focusing solely on low-quality data at any stage.
@@
We extract 70K samples from 7 domains within our training dataset, covering areas such as Arxiv, Law, Code, and Math. Using the 1.3B reference model, we calculate the PPL of all samples and present the PPL distributions in Figure 12. Significant differences are observed across various domains, with mean values ranging from 5 to 22 and varying degrees of variance. Data from domains such as Wikipedia typically exhibit lower PPL, whereas data from Reddit show higher PPL. This implies that sorting data by PPL could cause obvious shifts in domain representation between initial and later training stages, potentially exposing the model to overly homogeneous data during each training stage. Sachdeva et al. (2024) highlight that data diversity enhances model performance, and an uneven distribution may lead to a decline in performance.
@@
We extract 1 million samples from each of the four quadrant distributions and visualize them, as shown in Figure 8. We observe that each quadrant contains data from a wide range of sources, indicating that FRAME effectively ensures data diversity throughout the training process. Notably, there are distinct differences among the quadrants. For instance, $Q_{2}$ contains a large number of domain-specific datasets such as Wikipedia, and StackExchange. This aligns with common cognitive understanding, as these datasets may feature more complex sentence structures and terminology, which large models can learn effectively while small models may struggle to master, thus suitable for learning in the later stages of training. This also explains the excellent performance in knowledge-intensive tasks, as the texts in Wikipedia contain a wealth of factual knowledge. Additionally, the data distribution in $Q_{3}$ is relatively balanced, which is advantageous for the initial training phase. Presenting data from $Q_{3}$ at the beginning of the training process allows the model to learn from sufficiently diverse data, thereby establishing a strong foundation for subsequent learning stages.
@@
To assess distinct semantic properties within each quadrant’s data, we randomly select 1K samples from each quadrant and devise 10 language-text related traits for GPT-4o evaluation. For simplicity and precision, all traits are formulated as yes-or-no questions, and we calculate the percentage of samples meeting each trait’s criteria. Details of all traits are listed in Appendix C. As shown in Figure 9, the data across all four quadrants exhibit similar semantic properties. In addition, we use T5 Raffel et al. (2023) to obtain dense vectors of the samples and perform dimensionality reduction using t-SNE, as shown in Figure 10. We could observe that the data in the four quadrants do not have obvious distinctions at the semantic level. This indicates no obvious difference in the semantic distribution of data between the initial and later training stages, ensuring a diverse range of data is consistently encountered throughout the training process, and avoiding the collapse of the model into a certain preference, thereby damaging generalization performance.
Figure 9: Semantic analysis of different quadrants.
@@ 4 Related Works
Multi-stage pretraining has emerged as a pivotal strategy in the development of LLMs, enabling models to better capture and utilize diverse data characteristics by dividing the training process into distinct phasesPavlova (2025); Zhao et al. (2024); Liu et al. (2021b); Tan et al. (2022). Liu et al. (2021b) propose a multi-stage pretraining method that leverages various granularities of information, significantly boosting model performance. Tan et al. (2022) explore the use of multi-stage prompting to improve translation tasks, demonstrating its effectiveness in enhancing downstream applications.
In LLM pretraining, data preprocessing is key to ensuring dataset qualityDuan et al. (2025). Traditional methods use expert rules to filter low-quality data Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and remove duplicates Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). These approaches enhance quality but may lack semantic depth. To improve semantic selection, strategies involve using targeted sources or proxy models Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024). Classifiers automate selection, like the logistic regression model used by Du et al. (2022), with others employing complex scoring Zhang et al. (2024b); Sachdeva et al. (2024). QuRating Wettig et al. (2024) uses multiple raters for nuanced evaluation.
@@ 5 Conclusion
In this study, we propose the Four-quadRAnt Multi-stage prEtraining strategy (FRAME), a novel approach designed to boost the performance of LLMs by systematically organizing the pretraining process into four stages. Guided by the principle of achieving significant loss reductions four times, FRAME employs a strategic partitioning of pretraining data into four quadrants based on PPL and PD. The experimental results, showing a 16.8% average improvement across MMLU and CMMLU over random sampling, underscore the effectiveness of FRAME in optimizing pretraining data organization.
@@ Appendix A Ethical Considerations
Hyperparameter Value
Precision bfloat16
Layers 30
Hidden dimension 1920
Attention heads 15
Vocab size 131072
Sequence length 8192
Activation SiLU
Position embedding RoPE
Table 5: Model structure of 1.3B model
Hyperparameter Value
Precision bfloat16
Layers 32
Hidden dimension 2560
Attention heads 32
Vocab size 131072
Sequence length 8192
Activation SiLU
Position embedding RoPE
Table 6: Model structure of 3B model
@@ B.1 Experimental settings
Model configuration and training
We train reference models with 100M and 1.3B parameters on 500B tokens randomly sampled from the collected dataset, utilizing the Llama architecture Touvron et al. (2023). We use the 1.3B model to compute PPL and both the 100M and 1.3B models to calculate PD. For the main experiments, we train a 3B model. We set the batch size to 640 and the context window length to 8192. The initial learning rate is $2\times 10^{-4}$, with a warm-up phase of 375M tokens. We apply cosine learning rate scheduling with a weight decay of 0.1. We use the Adam optimizer to train the model within the Megatron framework Shoeybi et al. (2019). Furthermore, we validate our approach on a 1.3B model, using the same settings as the 3B model. We use Ascend 910B NPU for training. For the 3B model, we use 512 NPUs for training, each model takes over 180 hours; For the 1.3B models, we use 96 NPUs, each model takes 6 hours.
In Table 5 and 6, we present the model configuration of the 1.3B and 3B models.
Methods Steps MMLU CMMLU
Random 30K 25.4 25.9
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 26.7 26.2
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 25.3 25.5
Random 60K 25.9 25.8
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.5 25.5
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 33.4 35.4
Random 95K 24.8 25.6
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 26.0 26.0
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.6 42.6
Baselines
In the 3B model setting, we compare FRAME with several baselines:
  • Random: Data is organized in a completely random order, meaning there is no specific sequence or strategy for data input
  • PDPC Zhang et al. (2025): Utilizes PD as a model-aware curriculum learning indicator, organizing the data sequence through a preference function that aligns with the model’s inherent preferences.
  • $Q_{3}\rightarrow Q_{1}\rightarrow Q_{4}\rightarrow Q_{2}$: Adopts the same smoothing process as FRAME but prioritizes satisfying PD-related constraints to organize the data order.
In the 1.3B model setting, we also compare FRAME with other baselines:
  • Sequential: Organizes data by directly sorting according to PPL, PD, and QuRating Wettig et al. (2024), in either ascending or descending order.
  • Preference CL: Utilizes the same preference function as PDPC but replaces the indicators with PPL or QuRating. We adopt both S-shape and S-shape reverse function.
Methods Steps MMLU CMMLU
Random 30K 25.4 25.9
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.7 26.2
Random 60K 25.9 25.8
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.2 25.1
Random 95K 24.8 25.6
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.9 27.2
Figure 11: S-shape functions with a=10 and a=35.
Figure 12: PPL distribution across different sources.
Figure 13: PD Distribution across different sources.
@@ B.3 Benchmark Accuracy of Two-stage Pretraining Based on PD
Table 8 shows the model’s accuracy as training steps increase under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. The model consistently outperforms the Random model at most steps. Notably, at 95K steps, it exceeds the Random setting by 2.1% on MMLU and 1.6% on CMMLU, which validates PD as an effective metric. This leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
@@ B.5 Data Distribution
Figure 12 and 13 show the PPL and PD distribution of samples in our training set. The PPL of data from different domains shows significant variations, with mean values ranging from 5 to 22 and varying degrees of variance. Sorting data by PPL may lead to imbalanced domain representation between early and late training stages, exposing the model to overly homogeneous data at each stage and potentially degrading performance. In contrast, the distribution of PD is quite similar across different data sources, ensuring that each training phase includes diverse data, thereby maintaining pretraining efficiency and enhancing model performance.
@@
B.6 Additional evaluation results
In Figure 15, we present additional evaluation results on 3B models. It is evident that in the majority of benchmarks, FRAME significantly outperforms Random, underscoring the effectiveness of our proposed method.
(a) ARC-E
(b) ARC-C
(c) HellaSwag
(d) PIQA
(e) BBH
Figure 15: Performance of downstream tasks on 3B models with respect to training iterations. We compare FRAME with Random baseline.
B.7 Smoother LOSS Curve from FRAME